# How to avoid machine learning pitfalls: a guide for academic researchers

### v4

@@ Abstract
This document outlines some of the common mistakes that occur when using machine learning, and what can be done to avoid them. Whilst it should be accessible to anyone with a basic understanding of machine learning techniques, it was originally written for research students, and focuses on issues that are of particular concern within academic research, such as the need to do rigorous comparisons and reach valid conclusions. It covers five stages of the machine learning process: what to do before model building, how to reliably build models, how to robustly evaluate models, how to compare models fairly, and how to report results.
@@ 1 Introduction
It’s easy to make mistakes when applying machine learning (ML), and these mistakes can result in ML models that fail to work as expected when applied to data not seen during training and testing (Liao et al., 2021). This is a problem for practitioners, since it leads to the failure of ML projects. However, it is also a problem for society, since it erodes trust in the findings and products of ML (Gibney, 2022). This guide aims to help newcomers avoid some of these mistakes. It’s written by an academic, and focuses on lessons learnt whilst doing ML research in academia. Whilst primarily aimed at students and scientific researchers, it should be accessible to anyone getting started in ML, and only assumes a basic knowledge of ML techniques. However, unlike similar guides aimed at a more general audience, it includes topics that are of a particular concern to academia, such as the need to rigorously evaluate and compare models in order to get work published. To make it more readable, the guidance is written informally, in a Dos and Don’ts style. It’s not intended to be exhaustive, and references (with publicly-accessible URLs where available) are provided for further reading. Since it doesn’t cover issues specific to particular academic subjects, it’s recommended you also consult subject-specific guidance where available (e.g. Stevens et al. (2020) for medicine). Feedback is welcome, and it is expected that this document will evolve over time. For this reason, if you cite it, please include the arXiv version number (currently v4).
@@ 2.2 Do take the time to understand your data
Eventually you will want to publish your work. This is a lot easier to do if your data is from a reliable source, has been collected using a reliable methodology, and is of good quality. For instance, if you are using data collected from an internet resource, make sure you know where it came from. Is it described in a paper? If so, take a look at the paper; make sure it was published somewhere reputable, and check whether the authors mention any limitations of the data. Do not assume that, because a data set has been used by a number of papers, it is of good quality — sometimes data is used just because it is easy to get hold of, and some widely used data sets are known to have significant limitations (see Paullada et al. (2020) for a discussion of this). If you train your model using bad data, then you will most likely generate a bad model: a process known as garbage in garbage out. So, always begin by making sure your data makes sense. Do some exploratory data analysis (see Cox (2017) for suggestions). Look for missing or inconsistent records. It is much easier to do this now, before you train a model, rather than later, when you’re trying to explain to reviewers why you used bad data.
@@ 2.4 Do make sure you have enough data
If you don’t have enough data, then it may not be possible to train a model that generalises. Working out whether this is the case can be challenging, and may not be evident until you start building models: it all depends on the signal to noise ratio in the data set. If the signal is strong, then you can get away with less data; if it’s weak, then you need more data. If you can’t get more data — and this is a common issue in many research fields — then you can make better use of existing data by using cross-validation (see Do evaluate a model multiple times). You can also use data augmentation techniques (e.g. see Wong et al. (2016) and Shorten and Khoshgoftaar (2019); for time series data, see Iwana and Uchida (2021)), and these can be quite effective for boosting small data sets, though Don’t do data augmentation before splitting your data. Data augmentation is also useful in situations where you have limited data in certain parts of your data set, e.g. in classification problems where you have less samples in some classes than others, a situation known as class imbalance. See Haixiang et al. (2017) for a review of methods for dealing with this; also see Do choose metrics carefully. Another option for dealing with small data sets is to use transfer learning — see Do keep up with progress in deep learning (and its pitfalls). However, if you have limited data, then it’s likely that you will also have to limit the complexity of the ML models you use, since models with many parameters, like deep neural networks, can easily overfit small data sets (see Don’t assume deep learning will be the best approach). Either way, it’s important to identify this issue early on, and come up with a suitable strategy to mitigate against it.
@@ 2.7 Do think about how your model will be deployed
Why do you want to build an ML model? This is an important question, and the answer should influence the process you use to develop your model. Many academic studies are just that — studies — and not really intended to produce models that will be used in the real world. This is fair enough, since the process of building and analysing models can itself give very useful insights into a problem. However, for many academic studies, the eventual goal is to produce an ML model that can be deployed in a real world situation. If this is the case, then it’s worth thinking early on about how it is going to be deployed. For instance, if it’s going to be deployed in a resource-limited environment, such as a sensor or a robot, this may place limitations on the complexity of the model. If there are time constraints, e.g. a classification of a signal is required within milliseconds, then this also needs to be taken into account when selecting a model. Another consideration is how the model is going to be tied into the broader software system within which it is deployed; this procedure is often far from simple (see Sculley et al. (2015)). However, emerging approaches such as ML Ops aim to address some of the difficulties; see Tamburri (2020) for a review, and Shankar et al. (2022) for a discussion of common challenges when operationalising ML models.
@@ 3.1 Don’t allow test data to leak into the training process
It’s essential to have data that you can use to measure how well your model generalises. A common problem is allowing information about this data to leak into the configuration, training or selection of models (see Figure 1). When this happens, the data no longer provides a reliable measure of generality, and this is a common reason why published ML models often fail to generalise to real world data. There are a number of ways that information can leak from a test set. Some of these seem quite innocuous. For instance, during data preparation, using information about the means and ranges of variables within the whole data set to carry out variable scaling — in order to prevent information leakage, this kind of thing should only be done with the training data. Other common examples of information leakage are carrying out feature selection before partitioning the data (see Do be careful where and how you do feature selection), using the same test data to evaluate the generality of multiple models (see Do avoid sequential overfitting and Don’t always believe results from community benchmarks), and applying data augmentation before splitting off the test data (see Don’t do data augmentation before splitting your data). The best thing you can do to prevent these issues is to partition off a subset of your data right at the start of your project, and only use this independent test set once to measure the generality of a single model at the end of the project (see Do save some data to evaluate your final model instance). Be particularly careful if you’re working with time series data, since random splits of the data can easily cause leakage and overfitting — see Don’t ignore temporal dependencies in time series data for more on this. For a broader discussion of data leakage, see Kapoor and Narayanan (2022).
@@ 3.2 Do try out a range of different models
Generally speaking, there’s no such thing as a single best ML model. In fact, there’s a proof of this, in the form of the No Free Lunch theorem, which shows that no ML approach is any better than any other when considered over every possible problem (Wolpert, 2002). So, your job is to find the ML model that works well for your particular problem. There is some guidance on this. For example, you can consider the inductive biases of ML models; that is, the kind of relationships they are capable of modelling. For instance, linear models, such as linear regression and logistic regression, are a good choice if you know there are no important non-linear relationships between the features in your data, but a bad choice otherwise. Good quality research on closely related problems may also be able to point you towards models that work particularly well. However, a lot of the time you’re still left with quite a few choices, and the only way to work out which model is best is to try them all. Fortunately, modern ML libraries in Python (e.g. scikit-learn (Varoquaux et al., 2015)), R (e.g. caret (Kuhn, 2015)), Julia (e.g. MLJ (Blaom et al., 2020)) etc. allow you to try out multiple models with only small changes to your code, so there’s no reason not to try them all out and find out for yourself which one works best. However, Don’t use inappropriate models, and use a validation set, rather than the test set, to evaluate them (see Do avoid sequential overfitting). When comparing models, Do optimise your model’s hyperparameters and Do evaluate a model multiple times to make sure you’re giving them all a fair chance, and Do correct for multiple comparisons when you publish your results.
@@ 3.4 Do keep up with progress in deep learning (and its pitfalls)
Figure 2: See Do keep up with progress in deep learning (and its pitfalls). A rough history of neural networks and deep learning, showing what I consider to be the milestones in their development. For a far more thorough and accurate account of the field’s historical development, take a look at Schmidhuber (2015).
Whilst deep learning may not always be the best solution (see Don’t assume deep learning will be the best approach), if you are going to use deep learning, then it’s advisable to try and keep up with recent developments in this fast-moving field. Figure 2 summarises some of the important developments over time. Multilayer perceptrons (MLP) and recurrent neural networks (particularly LSTM) have been popular for some time, but are increasingly being replaced by newer models such as convolutional neural networks (CNN) (Li et al., 2021) and transformers (Lin et al., 2022). For example, transformers have become the go-to model for processing sequential data (e.g. natural language), and are increasingly being applied to other data types too, such as images (Khan et al., 2022). A prominent downside of both transformers and deep CNNs is that they have many parameters and therefore require a lot of data to train them. However, an option for small data sets is to use transfer learning, where a model is pre-trained on a large generic data set and then fine-tuned on the data set of interest (Han et al., 2021). Larger pre-trained models, many of which are freely shared on websites such as Hugging Face, are known as foundation models; see Zhou et al. (2023) for a survey. Whilst powerful, these come with their own set of pitfalls. For example, their ability to fully memorise input data is the cause of data security and privacy concerns (Li et al., 2023). The use of opaque, often poorly documented, training datasets also leads to pitfalls when fitting them into broader ML pipelines (see Do combine models (carefully) for more info) and comparing them fairly with other ML models (see Don’t assume a bigger number means a better model and Don’t always believe results from community benchmarks). For an extensive, yet accessible, guide to deep learning, see Zhang et al. (2021).
@@ 3.7 Do optimise your model’s hyperparameters
Many models have hyperparameters — that is, numbers or settings that affect the configuration of the model. Examples include the kernel function used in an SVM, the number of trees in a random forest, and the architecture of a neural network. Many of these hyperparameters significantly effect the performance of the model, and there is generally no one-size-fits-all. That is, they need to be fitted to your particular data set in order to get the most out of the model. Whilst it may be tempting to fiddle around with hyperparameters until you find something that works, this is not likely to be an optimal approach. It’s much better to use some kind of hyperparameter optimisation strategy, and this is much easier to justify when you write it up. Basic strategies include random search and grid search, but these don’t scale well to large numbers of hyperparameters or to models that are expensive to train, so it’s worth using tools that search for optimal configurations in a more intelligent manner. See Yang and Shami (2020) for a survey, and Bischl et al. (2021) for further guidance. It is also possible to use AutoML techniques to optimise both the choice of model and its hyperparameters, in addition to other parts of the data mining pipeline — see He et al. (2021) for a review.
@@ 3.8 Do avoid learning spurious correlations
Spurious correlations are features within data which are correlated with the target variable, but which have no semantic meaning. They are basically red herrings, and it’s not uncommon for ML models to pick up on them in training, and consequently fail to generalise well. A classic example is the tank problem. Legend [^2] There is some debate about whether this actually happened: see https://www.gwern.net/Tanks. has it that the US military were looking to train an ML model that could recognise tanks. However, because the tank pictures used in training were taken during different weather conditions to the non-tank pictures, the model ended up discriminating based on features such as the number of blue pixels in the sky, rather than the presence of a tank. More complex data tends to contain more of these spurious correlations, and more complex models have more capacity to overfit spurious correlations. This means that spurious correlations are a particular issue for deep learning, where approaches such as regularisation (see Do keep up with progress in deep learning (and its pitfalls)) and data augmentation (see Do make sure you have enough data) can help mitigate against this. However, spurious correlations can occur in all data sets and models, so it is always worth looking at your trained model to see whether it’s responding to appropriate features within your data — see Do look at your models.

### v5

@@ Abstract
Mistakes in machine learning practice are commonplace, and can result in a loss of confidence in the findings and products of machine learning. This guide outlines common mistakes that occur when using machine learning, and what can be done to avoid them. Whilst it should be accessible to anyone with a basic understanding of machine learning techniques, it focuses on issues that are of particular concern within academic research, such as the need to do rigorous comparisons and reach valid conclusions. It covers five stages of the machine learning process: what to do before model building, how to reliably build models, how to robustly evaluate models, how to compare models fairly, and how to report results.
@@ 1 Introduction
It’s easy to make mistakes when applying machine learning (ML), and these mistakes can result in ML models that fail to work as expected when applied to data not seen during training and testing (Liao et al., 2021). This is a problem for practitioners, since it leads to the failure of ML projects. However, it is also a problem for society, since it erodes trust in the findings and products of ML (Gibney, 2022).
This guide aims to help newcomers avoid some of these mistakes. It’s written by an academic, and focuses on lessons learnt whilst doing ML research in academia. Whilst primarily aimed at students and scientific researchers, it should be accessible to anyone getting started in ML, and only assumes a basic knowledge of ML techniques. However, unlike similar guides aimed at a more general audience, it includes topics that are of a particular concern to academia, such as the need to rigorously evaluate and compare models in order to get work published.
To make it more readable, the guidance is written informally, in a Dos and Don’ts style. It’s not intended to be exhaustive, and references (with publicly-accessible URLs where available) are provided for further reading. Since it doesn’t cover issues specific to particular academic subjects, it’s recommended that readers also consult subject-specific guidance where available, e.g. in clinical medicine (Stevens et al., 2020), genomics (Whalen et al., 2022), environmental research (Zhu et al., 2023), materials science (Karande et al., 2022), business and marketing (Van Giffen et al., 2022), computer security (Arp et al., 2022) and social science (Malik, 2020).
The review is divided into five sections. Before you start to build models covers issues that can occur early in the ML process, and focuses on the correct use of data and adequate consideration of the context in which ML is being applied. How to reliably build models then covers pitfalls that occur during the selection and training of models and their components. How to robustly evaluate models presents pitfalls that can lead to an incorrect understanding of model performance. How to compare models fairly then extends this to the situation where models are being compared, discussing how common pitfalls can lead to misleading findings. How to report your results focuses on reproducibility and factors that can lead to incomplete or deceptive reporting.
Changes
ML pitfalls are not static, and continue to evolve as ML develops. To address this, this guide has been updated annually since it was first released in 2021, and it will continue to be updated in the future. Feedback is welcome. If you cite it, please include the arXiv version number (currently v5222This version is published as “Avoiding machine learning pitfalls” in Patterns (Cell Press) (Lones, 2024)).
Changes from v4
Changes from v3
Changes from v2
Changes from v1
@@ 2.2 Do take the time to understand your data
Eventually you will want to publish your work. This is a lot easier to do if your data is from a reliable source, has been collected using a reliable methodology, and is of good quality. For instance, if you are using data collected from an internet resource, make sure you know where it came from. Is it described in a paper? If so, take a look at the paper; make sure it was published somewhere reputable, and check whether the authors mention any limitations of the data. Do not assume that, because a data set has been used by a number of papers, it is of good quality — sometimes data is used just because it is easy to get hold of, and some widely used data sets are known to have significant limitations (see Paullada et al. (2021) for a discussion of this). If you train your model using bad data, then you will most likely generate a bad model: a process known as garbage in garbage out. One way to avoid bad data sets is to build a direct relationship with people who generate data, since this increases the likelihood of obtaining a good-quality dataset that meets your needs. It also avoids problems of overfitting community benchmarks; see Don’t always believe results from community benchmarks. Yet regardless of where your data comes from, always begin by making sure that your data makes sense. Do some exploratory data analysis (see Cox (2017) for suggestions). Look for missing or inconsistent records. It is much easier to do this now, before you train a model, rather than later, when you’re trying to explain to reviewers why you used bad data.
@@ 2.3 Don’t look at all your data
2.4 Do clean your data
Even good-quality datasets will have issues. Some of these come from unavoidable noise or omissions in the data collection process, others are due to human error during collection or collation. Whatever the cause, it’s important to identify any issues, and do this before you start to build models. One common problem to look out for is data duplication, i.e. the unintentional inclusion of multiple copies of a data point. This can cause serious problems when a model is evaluated (see Don’t do data augmentation before splitting your data for an example), so should be identified and removed early on. Another common problem is missing values. Some models can cope with these, but many can’t, and so you’ll have to replace missing values with something else before they can be trained. There are various forms of imputation that can be used to achieve this; see Emmanuel et al. (2021) for a review. If you do imputation, be careful to avoid data leaks during imputation — see Don’t allow test data to leak into the training process. You should also check for outliers in your data, but only remove these if they are likely to be the result of noise or error rather than being natural extremes of the underlying data-generating process. For example, if a person’s age is greater than 150, then it’s probably an error; if it’s 110, then it could be a natural outlier. A related issue is meaningless or inconsistent data, for instance a person with a negative age. Data cleaning can be a time-consuming process, and becomes more challenging as the complexity of data increases. For this reason, many people have explored automating data cleaning using ML approaches; see Côté et al. (2024) for a review.
@@ 2.5 Do make sure you have enough data
If you don’t have enough data, then it may not be possible to train a model that generalises. Working out whether this is the case can be challenging, and may not be evident until you start building models: it all depends on the signal to noise ratio in the data set. If the signal is strong, then you can get away with less data; if it’s weak, then you need more data. If you can’t get more data — and this is a common issue in many research fields — then you can try using data augmentation techniques (see Wang et al. (2024), and for time series data, Iglesias et al. (2023)). These can be quite effective for boosting small data sets, though Don’t do data augmentation before splitting your data. Data augmentation is also useful in situations where you have limited data in certain parts of your data set, e.g. in classification problems where you have less samples in some classes than others, a situation known as class imbalance. See Haixiang et al. (2017) for a review of methods for dealing with this; also see Do choose metrics carefully. Another option for dealing with small data sets is to use transfer learning — see Do keep up with progress in deep learning (and its pitfalls). A danger when using small datasets is that different data partitions may be biased, for instance in terms of the quality or difficulty of data they contain. For this reason, it is advisable to consider frequent repartitioning. Cross-validation (see Do evaluate a model multiple times) is an efficient way of achieving this in small data sets. If you have limited data, then it’s also likely that you will have to limit the complexity of the ML models you use, since models with many parameters, like deep neural networks, can easily overfit small data sets (see Don’t assume deep learning will be the best approach). Regardless of how you handle the problem of limited data, it’s important to identify this issue early on, and come up with a suitable strategy to mitigate against it.
@@ 2.8 Do think about how your model will be deployed
Why do you want to build an ML model? This is an important question, and the answer should influence the process you use to develop your model. Many academic studies are just that — studies — and not really intended to produce models that will be used in the real world. This is fair enough, since the process of building and analysing models can itself give very useful insights into a problem. However, for many academic studies, the eventual goal is to produce an ML model that can be deployed in a real world situation. If this is the case, then it’s worth thinking early on about how it is going to be deployed. For instance, if it’s going to be deployed in a resource-limited environment, such as a sensor or a robot, this may place limitations on the complexity of the model. If there are time constraints, e.g. a classification of a signal is required within milliseconds, then this also needs to be taken into account when selecting a model. If using deep learning, then energy costs and carbon footprint may be a consideration, and if using LLMs, there may be further operational costs for hosting or accessing foundation models. Another consideration is how the model is going to be tied into the broader software system within which it is deployed; this procedure is often far from simple (see Sculley et al. (2015)). However, emerging approaches such as ML Ops aim to address some of the difficulties; see Kreuzberger et al. (2023) for a review, and Shankar et al. (2022) for a discussion of common challenges when operationalising ML models.
@@ 3.1 Don’t allow test data to leak into the training process
It’s essential to have data that you can use to measure how well your model generalises. A common problem is allowing information about this data to leak into the configuration, training or selection of models (see Figure 1). When this happens, the data no longer provides a reliable measure of generality, and this is a common reason why published ML models often fail to generalise to real world data. There are a number of ways that information can leak from a test set. Some of these seem quite innocuous. For instance, during data preparation, using information about the means and ranges of variables within the whole data set to carry out variable scaling or imputation — in order to prevent information leakage, these statistics should be calculated using only the training data. Other common examples of information leakage are carrying out feature selection before partitioning the data (see Do be careful where and how you do feature selection), using the same test data to evaluate the generality of multiple models (see Do avoid sequential overfitting and Don’t always believe results from community benchmarks), and applying data augmentation before splitting off the test data (see Don’t do data augmentation before splitting your data). The best thing you can do to prevent these issues is to partition off a subset of your data right at the start of your project, and only use this independent test set once to measure the generality of a single model at the end of the project (see Do save some data to evaluate your final model instance). There are also forms of data leakage which are specific to certain types of data. Time series data is particularly problematic, since the order of samples is significant, and random splits can easily cause leakage and overfitting — see Don’t ignore temporal dependencies in time series data for more on this. Even for non-time series data, the experimental conditions used to generate data sets may lead to temporal dependencies, or other problematic conditions such as duplicated or similar samples — see Do use an appropriate test set for an example. In order to prevent leakage, these kinds of issues need to be identified and taken into account when splitting data. For a broader discussion of data leakage, see Kapoor and Narayanan (2023).
@@ 3.2 Do try out a range of different models
Generally speaking, there’s no such thing as a single best ML model. In fact, there’s a proof of this, in the form of the No Free Lunch theorem, which shows that no ML approach is any better than any other when considered over every possible problem (Wolpert, 2002). So, your job is to find the ML model that works well for your particular problem. There is some guidance on this. For example, you can consider the inductive biases of ML models; that is, the kind of relationships they are capable of modelling. For instance, linear models, such as linear regression and logistic regression, are a good choice if you know there are no important non-linear relationships between the features in your data, but a bad choice otherwise. Good quality research on closely related problems may also be able to point you towards models that work particularly well. However, a lot of the time you’re still left with quite a few choices, and the only way to work out which model is best is to try them all. Fortunately, modern ML libraries, such as scikit-learn (Varoquaux et al., 2015) in Python, tidymodels (Kuhn and Wickham, 2020) in R, and MLJ (Blaom et al., 2020) in Julia, allow you to try out multiple models with only small changes to your code, so there’s no reason not to try them all out and find out for yourself which one works best. However, Don’t use inappropriate models, and use a validation set, rather than the test set, to evaluate them (see Do avoid sequential overfitting). When comparing models, Do optimise your model’s hyperparameters and Do evaluate a model multiple times to make sure you’re giving them all a fair chance, and Do correct for multiple comparisons when you publish your results.
@@ 3.4 Do keep up with progress in deep learning (and its pitfalls)
Figure 2: See Do keep up with progress in deep learning (and its pitfalls). A rough history of neural networks and deep learning, showing what I consider to be the milestones in their development. For a far more thorough account of the field’s historical development, take a look at Schmidhuber (2015).
Whilst deep learning may not always be the best solution (see Don’t assume deep learning will be the best approach), if you are going to use deep learning, then it’s advisable to try and keep up with recent developments in this fast-moving field. Figure 2 summarises some of the important developments over time. Multilayer perceptrons (MLP) and recurrent neural networks (particularly LSTM) have been around for some time, but have largely been subsumed by newer models such as convolutional neural networks (CNN) (Li et al., 2021) and transformers (Lin et al., 2022). For example, transformers have become the go-to model for processing sequential data (e.g. natural language), and are increasingly being applied to other data types too, such as images (Khan et al., 2022). A prominent downside of both transformers and deep CNNs is that they have many parameters and therefore require a lot of data to train them. However, an option for small data sets is to use transfer learning, where a model is pre-trained on a large generic data set and then fine-tuned on the data set of interest (Han et al., 2021). Larger pre-trained models, many of which are freely shared on websites such as Hugging Face, are known as foundation models; see Zhou et al. (2023) for a survey. Whilst powerful, these come with their own set of pitfalls. For example, their ability to fully memorise input data is the cause of data security and privacy concerns (Li et al., 2023). The use of opaque, often poorly documented, training datasets also leads to pitfalls when fitting them into broader ML pipelines (see Do combine models (carefully) for more info) and comparing them fairly with other ML models (see Don’t assume a bigger number means a better model and Don’t always believe results from community benchmarks). For an extensive, yet accessible, guide to deep learning, see Zhang et al. (2023).
@@ 3.7 Do optimise your model’s hyperparameters
Many models have hyperparameters — that is, numbers or settings that affect the configuration of the model. Examples include the kernel function used in an SVM, the number of trees in a random forest, and the architecture of a neural network. Many of these hyperparameters significantly effect the performance of the model, and there is generally no one-size-fits-all. That is, they need to be fitted to your particular data set in order to get the most out of the model. Whilst it may be tempting to fiddle around with hyperparameters until you find something that works, this is not likely to be an optimal approach. It’s much better to use some kind of hyperparameter optimisation strategy, and this is much easier to justify when you write it up. Basic strategies include random search and grid search, but these don’t scale well to large numbers of hyperparameters or to models that are expensive to train, so it’s worth using tools that search for optimal configurations in a more intelligent manner. See Bischl et al. (2023) for further guidance. It is also possible to use AutoML techniques to optimise both the choice of model and its hyperparameters, in addition to other parts of the machine learning pipeline — see Barbudo et al. (2023) for a review.
@@ 3.8 Do avoid learning spurious correlations
Figure 4: See Do avoid learning spurious correlations. The problem of spurious correlations in images, as illustrated by the tank problem. The images on the left are tanks, and those on the right are not tanks. However, the consistent background (blue for tanks, grey for others) means that these images can be classified by merely looking at the colours of pixels towards the top of the images, rather than having to recognise the objects in the images, resulting in a poor model.
Spurious correlations are features within data which are correlated with the target variable, but which have no semantic meaning. They are basically red herrings, and it’s not uncommon for ML models to pick up on them in training, and consequently fail to generalise well. A classic example is the tank problem. Legend [^3] There is some debate about whether this actually happened: see https://www.gwern.net/Tanks. has it that the US military were looking to train an ML model that could recognise tanks. However, because the tank pictures used in training were taken during different weather conditions to the non-tank pictures, the model ended up discriminating based on features such as the number of blue pixels in the sky, rather than the presence of a tank (see Figure 4 for an illustration). An ML model that uses such spurious correlations to perform classification would appear to be very good, in terms of its metric scores, but would not work in practice. More complex data tends to contain more of these spurious correlations, and more complex models have more capacity to overfit spurious correlations. This means that spurious correlations are a particular issue for deep learning, where approaches such as regularisation (see Do keep up with progress in deep learning (and its pitfalls)) and data augmentation (see Do make sure you have enough data) can help mitigate against this. However, spurious correlations can occur in all data sets and models, so it is always worth looking at your trained model to see whether it’s responding to appropriate features within your data — see Do look at your models.